BRICS Research Series, 2008

December 15, 2008

This document is also available as PostScript and DVI.

References

[RS-08-6] Abstract, PostScript, PDF, DVI.

Jacob Johannsen. An Investigation of Abadi and Cardelli’s Untyped Calculus of Objects. June 2008. xii+87 pp.

[RS-08-5] Abstract, PostScript, PDF, DVI.

Olivier Danvy and Jacob Johannsen. Inter-Deriving Semantic Artifacts for Object-Oriented Programming. June 2008. ii+13 pp. Extended version of a paper to appear in WoLLIC 2008.

[RS-08-4] Abstract, PostScript, PDF.

Olivier Danvy and Kevin Millikin. Refunctionalization at Work. June 2008. ii+25 pp. To appear in Science of Computer Programming. A preliminary version is available as the research report BRICS RS-07-7.

[RS-08-3] Abstract, PostScript, PDF, DVI.

Johan Munk. A Study of Syntactic and Semantic Artifacts and its Application to Lambda Definability, Strong Normalization, and Weak Normalization in the Presence of State. April 2008. xi+144 pp.

[RS-08-2] Abstract, PostScript, PDF, DVI.

Gudmund Skovbjerg Frandsen and Piotr Sankowski. Dynamic Normal Forms and Dynamic Characteristic Polynomial. April 2008. 21 pp. To appear in ICALP ’08.

[RS-08-1] Anders Møller. Static Analysis for Event-Based XML Processing. jan 2008. 23 pp. Appears in PLAN-X ’08.